OpenStudy (anonymous):

Part 2: (5 points) It’s your turn to be a game show host! As you know, in the game of Math Time, the contestants are given an answer and they must come up with the question that corresponds to the given answer. Your task for this portion of the assignment is to create two different “answers” (and the questions that accompany them) that the host could use for the final round of Math Time.

OpenStudy (anonymous):

A frustum of a regular hexagonal pyramid has an upper base edge of 16 ft. and a lower base edge of 28 ft. If the lateral area of the frustum is 1,716 ft.2, find the altitude of the frustum. Atrapezoid = 1,716/6 Atrapezoid = 286 ft.2 A = ½ (a + b)h 286 = ½ (16 + 28)h h = 286/½(16+28) h = 13 ft. l = √h2 + (r2 – r1)2 l2 = 132 + (14 – 8)2 l2 = 169 + 36 l = √205 l = √h2 + (r2 – r1)2 205 = h2 + (28 – 16)2 205 = h2 + 144 h = √61 ft. ------------ HAHAHA. :))

OpenStudy (anonymous):

omg ur so smart at math ..thanks btw

OpenStudy (anonymous):

That's one of my homework. Hahaha.

OpenStudy (anonymous):

thnks !]